# A Novel Reinforcement Learning Model for Post-Incident Malware Investigations

### v1

@@ Abstract
This Research proposes a Novel Reinforcement Learning (RL) model to optimise malware forensics investigation during cyber incident response. It aims to improve forensic investigation efficiency by reducing false negatives and adapting current practices to evolving malware signatures. The proposed RL framework leverages techniques such as Q-learning and the Markov Decision Process (MDP) to train the system to identify malware patterns in live memory dumps, thereby automating forensic tasks. The RL model is based on a detailed malware workflow diagram that guides the analysis of malware artefacts using static and behavioural techniques as well as machine learning algorithms. Furthermore, it seeks to address challenges in the UK justice system by ensuring the accuracy of forensic evidence. We conduct testing and evaluation in controlled environments, using datasets created with Windows operating systems to simulate malware infections. The experimental results demonstrate that RL improves malware detection rates compared to conventional methods, with the RL model’s performance varying depending on the complexity and learning rate of the environment. The study concludes that while RL offers promising potential for automating malware forensics, its efficacy across diverse malware types requires ongoing refinement of reward systems and feature extraction methods.
@@ I Introduction
Reinforcement learning, on the other hand, enables models to generate adversarial malware that can bypass detection by portable executable (PE) malware classifiers. Quertier et al., [8], note that the RL framework "Gym-Malware" achieves an evasion rate of up to 16%, while the RL framework "MAB-malware" can achieve an evasion rate of over 75% in a black-box setting. Furthermore, according to Quertier et al., [8], adversarial knowledge defines attacks as either "white box," where the adversary has complete access to the model, or "black box," where the adversary has no knowledge of the model and can only obtain classification results through a limited number of attempts [11]. These adversarial samples, which are falsely classified as benign, can improve the detection accuracy of malware detectors by 16% to 94% [2]. Reinforcement learning’s ability to automate malware evasion shows how detailed sequences of adversarial actions can train antivirus and malware detection systems, enhancing their effectiveness in combating malware [8].
@@ II-B Adversarial RL for Malware Detection
Ebrahimi et al., [6] research improves cybersecurity by addressing adversarial attacks on ML-based malware detection. Specifically, their innovative use of adversarial reinforcement learning (RL) enhances malware detector robustness in dynamic environments. The method improves dynamic interactions between the detector and the adversary by including adversarial agents that create false samples. However, reliance on MalVAC and lack of comparative assessment limit the research’s generalizability. Furthermore, more theoretical exploration of convergence properties could improve the framework. Despite these flaws, their focus on future work like optimisation solvers and reducing false positives shows forward-thinking.
@@ III-A Experimental Setup and Dataset Generation
To implement and validate the proposed reinforcement learning malware investigation framework, a comprehensive malware dataset was created using the London Metropolitan University Digital Forensics Laboratory [15]. We set up thirteen virtual machines in a secure network to prevent unintended malware spread. We uploaded different ISO files of the Windows operating system to ensure a diverse test environment. We introduced malware to each virtual machine, took snapshots of infected and uninfected states, and produced 26 RAM files. To uncover malware behaviours, we analysed these files using the Volatility framework. Finally, we created a detailed workflow diagram to facilitate the training and validation of the model.
@@ III-B Malware Workflow Diagram Creation
The research methodology covers the experimental setup, dataset generation, and development of a malware analysis workflow. This workflow is central to the reinforcement learning framework, integrating techniques such as data collection, examination, and analysis [16]. We analysed live memory dumps from 13 versions of Windows, both infected and uninfected, using the Volatility framework to detect anomalies and malware artefacts. The analysis phase employs static, signature-based, behavioural techniques, and machine learning algorithms. The workflow diagram maps typical malware behaviours and improves post-incident forensic investigations, supporting the reinforcement learning model’s training and validation.
@@ III-D The Unified Markov Decision Process
III-E The Proposed RL Post-Incident Malware Investigation Framework
The six main parts of the Reinforcement Learning Post-Incident Malware Investigation Framework are shown in Figure 2. These components include data collection, workflow diagram mapping, MDP model implementation, environmental dependencies, the MDP solver, and continuous learning and adaptation. Initially, data collection involves acquiring live memory dumps from Windows operating systems. Following that, data analysis focusses on analysing the collected data to identify anomalies, compromise indicators, and potential malware rtifacts. The workflow diagram outlines a comprehensive approach to identifying malware infections using static analysis, signature-based analysis, behavioural analysis, and machine learning algorithms. To extract features from the identified processes, we use the AWK module. Additionally, listing DLLs is essential for tracking loaded DLLs for each process. Furthermore, monitoring open handles is crucial for keeping track of the open handles associated with each process. Moreover, collecting network data ensures the acquisition of all pertinent network-related information. Registry hive analysis involves identifying registry hives and listing their keys. We duplicate processes into executable files and check them against known malware databases to determine their maliciousness or benignity. Additionally, addressable memory is duplicated to conduct grep searches using specific keywords. The state spaces are designed to align with the malware workflow diagram, encompassing 67 unique states. Based on this workflow, actions are defined, ranging from three to ten, exposing the agent to 109 distinct actions within a defined environment. We create the MDP solver by formulating the unified Markov Decision Process (MDP) models, which we base on the subsections of the proposed RL Post-Incident Malware Investigation Model. Finally, we divide the setup environment dependencies into three sections: creating dependencies and gym environments, importing required libraries, and implementing the training data for continuous learning and adaptation.
Figure 2: The Proposed RL Post-Incident Malware Investigation Framework
@@ VII Conclusion
This paper presents a novel reinforcement learning (RL) model and framework for post-incident malware forensics investigations, designed to surpass the capabilities of human forensic experts. The model accelerates malware analysis and identifies both known and unknown threats by integrating various techniques, such as live memory dumps from Windows systems. A unified Markov Decision Process (MDP) framework was developed, featuring three environments: BlankEnvironment, BlankEnvironment_with_Rewards, and BlankEnvironment_with_Time, each with distinct reward mechanisms to optimise malware analysis. The RL agent, utilising Q-learning and epsilon-greedy exploration, iteratively refines its policy and decision-making process, improving malware identification accuracy. Experimental tests, including simulations with malware like WannaCry and Cerber, demonstrated that performance depends on learning rates and environment complexity. The study focusses on hyperparameter tuning and continuous learning to improve the performance of RL models. It shows that reward systems, feature extraction, and hybrid analysis could all use more optimisation.

### v2

@@ Abstract
This Research proposes a Novel Reinforcement Learning (RL) model to optimise malware forensics investigation during cyber incident response. It aims to improve forensic investigation efficiency by reducing false negatives and adapting current practices to evolving malware signatures. The proposed RL System leverages techniques such as Q-learning and the Markov Decision Process (MDP) to train the system to identify malware patterns in live memory dumps, thereby automating forensic tasks. The RL model is based on a detailed malware workflow diagram that guides the analysis of malware artefacts using static and behavioural techniques as well as machine learning algorithms. Furthermore, it seeks to address challenges in the UK justice system by ensuring the accuracy of forensic evidence. We conduct testing and evaluation in controlled environments, using datasets created with Windows operating systems to simulate malware infections. The experimental results demonstrate that RL improves malware detection rates compared to conventional methods, with the RL model’s performance varying depending on the complexity and learning rate of the environment. The study concludes that while RL offers promising potential for automating malware forensics, its efficacy across diverse malware types requires ongoing refinement of reward systems and feature extraction methods.
@@ I Introduction
Reinforcement learning, on the other hand, enables models to generate adversarial malware that can bypass detection by portable executable (PE) malware classifiers. Quertier et al., [8], note that the RL System "Gym-Malware" achieves an evasion rate of up to 16%, while the RL System "MAB-malware" can achieve an evasion rate of over 75% in a black-box setting. Furthermore, according to Quertier et al., [8], adversarial knowledge defines attacks as either "white box," where the adversary has complete access to the model, or "black box," where the adversary has no knowledge of the model and can only obtain classification results through a limited number of attempts [11]. These adversarial samples, which are falsely classified as benign, can improve the detection accuracy of malware detectors by 16% to 94% [2]. Reinforcement learning’s ability to automate malware evasion shows how detailed sequences of adversarial actions can train antivirus and malware detection systems, enhancing their effectiveness in combating malware [8].
@@ II-B Adversarial RL for Malware Detection
Ebrahimi et al., [6] research improves cybersecurity by addressing adversarial attacks on ML-based malware detection. Specifically, their innovative use of adversarial reinforcement learning (RL) enhances malware detector robustness in dynamic environments. The method improves dynamic interactions between the detector and the adversary by including adversarial agents that create false samples. However, reliance on MalVAC and lack of comparative assessment limit the research’s generalizability. Furthermore, more theoretical exploration of convergence properties could improve the System. Despite these flaws, their focus on future work like optimisation solvers and reducing false positives shows forward-thinking.
@@ III-A Experimental Setup and Dataset Generation
To implement and validate the proposed reinforcement learning malware investigation System, a comprehensive malware dataset was created using the London Metropolitan University Digital Forensics Laboratory [15]. We set up thirteen virtual machines in a secure network to prevent unintended malware spread. We uploaded different ISO files of the Windows operating system to ensure a diverse test environment. We introduced malware to each virtual machine, took snapshots of infected and uninfected states, and produced 26 RAM files. To uncover malware behaviours, we analysed these files using the Volatility System. Finally, we created a detailed workflow diagram to facilitate the training and validation of the model.
@@ III-B Malware Workflow Diagram Creation
The research methodology covers the experimental setup, dataset generation, and development of a malware analysis workflow. This workflow is central to the reinforcement learning System, integrating techniques such as data collection, examination, and analysis [16]. We analysed live memory dumps from 13 versions of Windows, both infected and uninfected, using the Volatility System to detect anomalies and malware artefacts. The analysis phase employs static, signature-based, behavioural techniques, and machine learning algorithms. The workflow diagram maps typical malware behaviours and improves post-incident forensic investigations, supporting the reinforcement learning model’s training and validation.
@@ III-D The Unified Markov Decision Process
III-E The Proposed RL Post-Incident Malware Investigation System
The six main parts of the Reinforcement Learning Post-Incident Malware Investigation System are shown in Figure 2. These components include data collection, workflow diagram mapping, MDP model implementation, environmental dependencies, the MDP solver, and continuous learning and adaptation. Initially, data collection involves acquiring live memory dumps from Windows operating systems. Following that, data analysis focusses on analysing the collected data to identify anomalies, compromise indicators, and potential malware rtifacts. The workflow diagram outlines a comprehensive approach to identifying malware infections using static analysis, signature-based analysis, behavioural analysis, and machine learning algorithms. To extract features from the identified processes, we use the AWK module. Additionally, listing DLLs is essential for tracking loaded DLLs for each process. Furthermore, monitoring open handles is crucial for keeping track of the open handles associated with each process. Moreover, collecting network data ensures the acquisition of all pertinent network-related information. Registry hive analysis involves identifying registry hives and listing their keys. We duplicate processes into executable files and check them against known malware databases to determine their maliciousness or benignity. Additionally, addressable memory is duplicated to conduct grep searches using specific keywords. The state spaces are designed to align with the malware workflow diagram, encompassing 67 unique states. Based on this workflow, actions are defined, ranging from three to ten, exposing the agent to 109 distinct actions within a defined environment. We create the MDP solver by formulating the unified Markov Decision Process (MDP) models, which we base on the subsections of the proposed RL Post-Incident Malware Investigation Model. Finally, we divide the setup environment dependencies into three sections: creating dependencies and gym environments, importing required libraries, and implementing the training data for continuous learning and adaptation.
Figure 2: The Proposed RL Post-Incident Malware Investigation System
@@ VII Conclusion
This paper presents a novel reinforcement learning (RL) model and System for post-incident malware forensics investigations, designed to surpass the capabilities of human forensic experts. The model accelerates malware analysis and identifies both known and unknown threats by integrating various techniques, such as live memory dumps from Windows systems. A unified Markov Decision Process (MDP) System was developed, featuring three environments: BlankEnvironment, BlankEnvironment_with_Rewards, and BlankEnvironment_with_Time, each with distinct reward mechanisms to optimise malware analysis. The RL agent, utilising Q-learning and epsilon-greedy exploration, iteratively refines its policy and decision-making process, improving malware identification accuracy. Experimental tests, including simulations with malware like WannaCry and Cerber, demonstrated that performance depends on learning rates and environment complexity. The study focuses on hyperparameter tuning and continuous learning to improve the performance of RL models. It shows that reward systems, feature extraction, and hybrid analysis could all use more optimisation. In future work, we intend to integrate a rule-based expert system (RBES) to capture expertise generated by our system, generalise it and use it directly in future investigations where the case investigated involves similar machine architecture and configurations, this will certainly improve the performance and maintain consistency [18].